100 mL of 0.10M HNO2 is titrated with a 0.10M NaOH solution. find the pH at the equivalence point
1
Expert's answer
2017-04-21T12:25:06-0400
Chemical equation:
HNO2 + NaOH → NaNO2 + H2O
The volume of NaOH at the equivalence point equals:
V(NaOH) = (C(HNO2)∙V(HNO2))/(C(NaOH)) = (0.10∙0.100)/0.10 = 0.100 (L)
At the equivalence point we have buffer system. This buffer solution is a mixture of a weak acid HNO2 and a salt of this acid – NaNO2.
The pH value of acidic buffer system we can calculate, using Henderson - Hasselbalch equation:
pH = pKa + lg(C(NaNO2))/(C(HNO2)), where pKa = - lg Ka (Ka is acid constant dissociation).
Ka(HNO2) = 4.0 · 10-4
We can determine the pKa value:
pKa = - lg 4.0 · 10-4 = 3.4
The number of moles of acid and sodium hydroxide are:
n = C·V = 0.10·0.100 = 0.01 (moles)
So, plug in values to find
pH = 3.4 + lg(0.01 )/0.01 = 3.4

Answer: pH = 3.4.